IBM 1401/1440/1460/1410/7010 Character Code Chart in BCD Order

PRINTS  IBM DEFINED  OP  CARD BCD  PRINTS  IBM DEFINED  OP  CARD BCD
AS  ASCII  CHARACTER  CODE CODE CODE AS  ASCII CHARACTER  CODE CODE CODE
BLANK BLANK  C  -  -  -  11  B
1  1  1  R  1  1  J  J  J  B¶  11-1  CB1
2  2  2  W  2  2  K  K  K  SS  11-2  CB2
3  3  3  WR  3  C21  L  L  L  LCA  11-3  B21
4  4  4  P  4  4  M  M  M  MCW  11-4  CB4
5  5  5  RP  5  C41  N  N  N  NOP  11-5  B41
6  6  6  WP  6  C42  O  O  O  11-6  B42
7  7  7  WRP  7  421  P  P  P  MCM* 11-7  CB421
8  8  8  SRF* 8  8  Q  Q  Q  SAR‡  11-8  CB8
9  9  9  SPF*  9  C81  R  R  R  ¶  11-9  B81
0  0  0  0  C82  -  !  ! (Minus Zero)  ZS  11-0  B82
#=  ##=  #=  MA†  3-8  821  $  $  $  11-3-8 CB821
@ ’  @@ ’  @ ’  M*  4-8  C84  *  *  *  11-4-8 B84
:  :  5-8  841  ]  )  11-5-8 CB841
>  >  6-8  842  ;  ;  11-6-8 CB842
( { {  √  7-8  C8421  Δ (Mode Change)  11-7-8 B8421
= |  ˆ  ¢(/ b on 1410)  2-8¶♭  A  &+  &&+ &+  12  CBA
/  /  /  CS  0-1  CA1  A  A  A  A  12-1  BA1
S  S  S  S  0-2  CA2  B  B  B  B  12-2  BA2
T  T  T  T§¶  0-3  A21  C  C  C  C  12-3  CBA21
U  U  U  CU  0-4  CA4  D  D  D  MN  12-4  BA4
V  V  V  BWZ  0-5  A41  E  E  E  MCE  12-5  CBA41
W  W  W  BBE* 0-6  A42  F  F  F  CC  12-6  CBA42
X  X  X  *¶  0-7  CA421 G  G  G  S?R¶  12-7  BA421
Y  Y  Y  MZ  0-8  CA8  H  H  H  SBR‡  12-8  BA8
Z  Z  Z  MCS  0-9  A81  I  I  I  12-9  CBA81
= |  ’ | |  = | Record Mark  0-2-8  A82  &+  ?  ? (Plus Zero)  ZA  12-0  CBA82
,  ,  ,  SW  0-3-8  CA821 .  .  .  H  12-3-8 BA821
% (  %%  % (  D*  0-4-8  A84  □)  )  □)  CW  12-4-8 CBA84
=∼∼  = Word separator  0-5-8  CA841  [  (  12-5-8 BA841
\  ’ Apostrophe  0-6-8  CA842  <  <  12-6-8 BA842
+””  ” Tape Segment  0-7-8  A8421  ” } } ≡ | Group Mark  12-7-8 CBA8421

Where two characters are shown for printing, the first is printed by the A chain, and the second by the H chain. Where three characters are shown for ASCII, they are traditional SimH encoding, Pierce’s primary encoding, and Pierce’s alternative encoding. The IBM defined character encoding is from Figure 267 on page 170 of A24-1403, Reference Manual: IBM 1401 Data Processing System. † If more than 4k memory on 1401, ‡ Store address register feature, * Other special feature, §1460, ¶1410, ♭ 1440. D, T and X have different meanings in 1401, 1460 and 1410. 1410 does not interpret #, H, P, Q, Y or 1-9. G, J and R are 1410 only. M and L on 1410 are only for I/O. 1440 does not interpret 1-9. P, #, % and @ are standard on 1460. % and @ are standard on 1410. 7010 and 1410 are compatible. ¶♭ Without no-charge RPQ 898148, the A-bit coding must be program generated in the 1401 (it cannot be read from a card; it punches as zero). It is used in conjunction with the parity track to indicate a blank position on tape written using even parity. The Word Separator (CA841) is used on tape to indicate that the next character has a word mark.

IBM 1401/1440/1460/1410/7010 Character Code Chart in Collating Sequence

PRINTS  IBM DEFINED  OP  CARD BCD  PRINTS  IBM DEFINED OP  CARD BCD
AS  ASCII  CHARACTER  CODE CODE CODE  AS  ASCII CHARACTER  CODE CODE CODE
BLANK BLANK  C  G  G  G  S?R¶  12-7  BA421
.  .  .  H  12-3-8 BA821  H  H  H  SBR‡  12-8  BA8
□)  )  □)  CW  12-4-8 CBA84  I  I  I  12-9  CBA81
[  (  12-5-8 BA841  -  !  ! (Minus Zero) ZS  11-0  B82
<  <  12-6-8 BA842  J  J  J  B¶  11-1  CB1
” } }  ≡ | Group Mark  12-7-8 CBA8421 K  K  K  SS  11-2  CB2
&+  &&+  &+  12  CBA  L  L  L  LCA  11-3  B21
$  $  $  11-3-8 CB821  M  M  M  MCW  11-4  CB4
*  *  *  11-4-8 B84  N  N  N  NOP  11-5  B41
]  )  11-5-8 CB841  O  O  O  11-6  B42
;  ;  11-6-8 CB842  P  P  P  MCM* 11-7  CB421
Δ (Mode Change)  11-7-8 B8421  Q  Q  Q  SAR‡  11-8  CB8
-  -  -  11  B  R  R  R  ¶  11-9  B81
/  /  /  CS  0-1  CA1  = |  ’ | |  = | Record Mark  0-2-8  A82
,  ,  ,  SW  0-3-8  CA821  S  S  S  S  0-2  CA2
% (  %% (  % (  D*  0-4-8  A84  T  T  T  T§¶  0-3  A21
=∼∼  = Word separator  0-5-8  CA841  U  U  U  CU  0-4  CA4
\  ’ Apostrophe  0-6-8  CA842  V  V  V  BWZ  0-5  A41
+””  ” Tape Segment  0-7-8  A8421  W  W  W  BBE*  0-6  A42
= |  ˆ  ¢(/ b on 1410)  2-8¶♭  A  X  X  X  *¶  0-7  CA421
#=  ##=  #=  MA†  3-8  821  Y  Y  Y  MZ  0-8  CA8
@ ’  @@ ’  @ ’  M*  4-8  C84  Z  Z  Z  MCS  0-9  A81
:  :  5-8  841  0  0  0  0  C82
>  >  6-8  842  1  1  1  R  1  1
( { {  √  7-8  C8421  2  2  2  W  2  2
&+  ?  ? (Plus Zero)  ZA  12-0  CBA82  3  3  3  WR  3  C21
A  A  A  A  12-1  BA1  4  4  4  P  4  4
B  B  B  B  12-2  BA2  5  5  5  RP  5  C41
C  C  C  C  12-3  CBA21  6  6  6  WP  6  C42
D  D  D  MN  12-4  BA4  7  7  7  WRP  7  421
E  E  E  MCE  12-5  CBA41  8  8  8  SRF*  8  8
F  F  F  CC  12-6  CBA42  9  9  9  SPF*  9  C81

Where two characters are shown for printing, the first is printed by the A chain, and the second by the H chain. Where three characters are shown for ASCII, they are traditional SimH encoding, Pierce’s primary encoding, and Pierce’s alternative encoding. The IBM defined character encoding is from Figure 267 on page 170 of A24-1403, Reference Manual: IBM 1401 Data Processing System. † If more than 4k memory on 1401, ‡ Store address register feature, * Other special feature, §1460, ¶1410, ♭ 1440. D, T and X have different meanings in 1401, 1460 and 1410. 1410 does not interpret #, H, P, Q, Y or 1-9. G, J and R are 1410 only. M and L on 1410 are only for I/O. 1440 does not interpret 1-9. P, #, % and @ are standard on 1460. % and @ are standard on 1410. 7010 and 1410 are compatible. ¶♭ Without no-charge RPQ 898148, the A-bit coding must be program generated in the 1401 (it cannot be read from a card; it punches as zero). It is used in conjunction with the parity track to indicate a blank position on tape written using even parity. The Word Separator (CA841) is used on tape to indicate that the next character has a word mark.